OpenStudy (anonymous):

2x/(4x+2) x (14x+7)/6 and 5/(x-2) - 3/(x-2). I got both of these problems wrong on my quiz and I can't understand why. For the second problem, I assume 2/(x-2) would be the simplest answer but that wasn't one of the choices. Can anybody explain these problems to me?

OpenStudy (campbell_st):

do some factoring 1st \[\frac{2x}{2(2x + 1)} \times \frac{7(2x +1)}{6} \] you may see you can cancel 2 in the numerator and denominator of the 1st fraction and (2x + 1) will cancel leaving the answer \[\frac{7x}{6}\]

OpenStudy (campbell_st):

the 2nd question, both fractions have the same denominator and you can factor the denominator. so you get \[\frac{5}{(2x -2)} - \frac{3}{(2x - 2)} = \frac{5 - 3}{2(x -1)}\] you just need to simplify it by cancelling a common factor.

OpenStudy (anonymous):

Wow, the first one so simple. I feel stupid. The second one, though, is not one of the answer choices. They're listed here: http://oi44.tinypic.com/8xm1ea.jpg

OpenStudy (anonymous):

for the 2nd question, is one of the denominators x+2?? if so, which one is it??

OpenStudy (anonymous):

Not unless the teacher made a typo, which I possible.

OpenStudy (anonymous):

i think so.........